OpenStudy (anonymous):

I need help finding the axis of symmetry. f(x)= −2(x − 4)2 + 2

OpenStudy (anonymous):

@E.ali

OpenStudy (anonymous):

is that -2(x-4)^(2) + 2

OpenStudy (anonymous):

Yes, sorry about that f(x)= −2(x − 4)^2 + 2

OpenStudy (amistre64):

given a vertex: (p,q) the axis of symmetry is defined as x=p you have a vertex form of a quadratic .... can you spot the x part?

OpenStudy (amistre64):

if you expand it all out to the ax^2 + bx + c form .... the axis can be defined as: -b/(2a)

OpenStudy (anonymous):

2x-8^2 + 2 x=2 -b/(2a) = -8/(2)2 -8/(2)2 = -8/4 -8/4=-2

OpenStudy (amistre64):

i think your calcs are in error

OpenStudy (amistre64):

square first, then distribute

OpenStudy (amistre64):

−2(x − 4)^2 + 2 −2(x^2-8x +16) + 2 etc ...

OpenStudy (anonymous):

-b/(2a) = -16/(2)2 -16/(2)2 = -16/4 -16/4=-4?

OpenStudy (anonymous):

−2(x^2-8x +16) + 2 -2x^2-16x-32+2?

OpenStudy (anonymous):

-2x^2-16x-30 -4x-16x-30 -20x-30?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

-2x^2-16x-30 this is good, now -b = --16 2a = 2(-2) what does that simplify to?

OpenStudy (anonymous):

-b=16 2a=-4

OpenStudy (amistre64):

doh!! -2(-8) = 16 .... not -16

OpenStudy (amistre64):

-16/-4 = 4

OpenStudy (amistre64):

-2x^2 +16x -30 this is good, now -b = -16 2a = 2(-2) -16 ---- = 4, thats the axis -4

OpenStudy (anonymous):

Thanks, could you help me find the axis of symmetry for g(x) = 5x2 − 10x + 7 as well?